OpenStudy (anonymous):

Determine whether the lines -x-2y = 3 and (x+2y)/2 = 1 are parallel, perpendicular, or neither. What to do with the (x+2y)/2?

OpenStudy (anonymous):

Distribute the (1/2) leaving you with: (1/2)x+y=1 then subtract the (1/2)x giving: y=(-1/2)x+1 Solve the other: -2y=3+x y=(-3/2)+(-1/2)x So, since the slopes are the same, they are parallel. :)

OpenStudy (anonymous):

Did that make sense?

OpenStudy (anonymous):

How do you distribute (1/2)?

OpenStudy (anonymous):

I've only used the distributive property for whole numbers..

OpenStudy (anonymous):

Do you agree that dividing by 2 is the same as multiplying by (1/2)? You don't treat the distributive property any different for fractions (or anything for that matter :P).

OpenStudy (anonymous):

For example: Say you have x(6+3) you add the 6 and 3 right? Or x(9)=9x. But check this out. Instead, distribute the x. Giving: 6x+3x. But that also equals 9x. Don't let the fraction scare you, it distributes just as any whole number :)

OpenStudy (anonymous):

So I just imagine a 1 above the dash, over the 2?

OpenStudy (anonymous):

Yes. For example: Take 9/3. That equals 3 right? Well that is the same as 9(1/3).

OpenStudy (anonymous):

Ok. Thanks.

OpenStudy (anonymous):

No problem, if you run into any problems just post the questions and I can help you (or someone else if you wish)